OpenStudy (anonymous):

The function arccot is defined as the inverse of the cotangent function restricted to the interval (0,π). Suppose we define a function f by f(x)=cot(x) for x in the interval (2⋅π,3⋅π) and let g be the inverse function to f. find g(-1)

OpenStudy (anonymous):

Hint: Substitute x with -1 for g(x) and evaluate. Do the same for g(√3).

OpenStudy (anonymous):

Suppose we define a function f by f(x)=cot(x) for x in the interval (2*π,3*π) and let g be the inverse function to f. Give the domain of g. Find the values of g⁡(sqrt(3))= and g(⁡-1)=

OpenStudy (anonymous):

i found g(sqrt(3)) to be 13/6 pi but i cant find g(-1)

OpenStudy (anonymous):

Substitute it

OpenStudy (anonymous):

what is g(x) look like

OpenStudy (ash2326):

\[f(x)=cot x\] \[g(x)= 2\pi+ \cot^{-1} x\] \[g(-1)=2pi+ \cot^{-1} (-1)\] arccot -1= 3pi/4 so \[g(-1)= 2\pi+3\pi/4\]

OpenStudy (anonymous):

thanks so much

OpenStudy (dumbcow):

y = f(x) = 1/tan(x) --> tan(x) = 1/y --> x = g(y) = arctan(1/y) g(-1) = arctan(-1) = 3pi/4 in interval 2pi to 3pi, add 2pi 2pi + 3pi/4 = 11pi/4

OpenStudy (anonymous):

ok i got it